# spacing=nonfrench

### 04 Mar 2026 v1

@@ 1.2. Torsion points
Whenever we talk about the roots or the fixed points of a power series, we mean its roots or its fixed points, respectively, in the $p$-adic open unit disk $\mathfrak{m}_{\mathbb{C}_{p}}\!$ of $\mathbb{C}_{p}$. A noninvertible formal power series $s_{1}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s_{1}^{\prime}(0)$ is nonzero can have no other fixed points than 0, but many roots. The set of roots of $s_{1}$ and its iterates is denoted by
$\Lambda(s_{1})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid\exists\,n\geqslant 0,\;s_{1}^{\circ n}(x)=0\}.$
By contrast, a nontorsion invertible power series $s_{2}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ can have no other roots than 0, but many fixed points. When $s_{1}$ and $s_{2}$ commute with each other, Lubin showed in [Lub94, Proposition 3.2] that $\Lambda(s_{1})$ is the set of periodic points of $s_{2}$. Knowing more about these points (their valuation, how many there are for a given valuation) is a key point in our arguments. This can be achieved by means of the Weierstrass preparation theorem and Newton polygons. The use of Newton polygons requires fixing a valuation; for this, we always use the valuation $\upsilon_{K}$ normalized on $K$ so that $\upsilon_{K}(p)=e$.
@@
We proceed by induction on $n$. The base case ($n=1$) follows directly from the hypotheses on $f$ and [Li96, Theorem 4.1]. The decreasing part of the Newton polygon of $f$ consists of a single line of slope e/(p1)-e/(p-1), implying that the Weierstrass polynomial associated to $f(X)/X$ is irreducible over $K$ since (e,p1)=1(e,p-1)=1. Assume now that the result holds for $n\geqslant 1$. By multiplicativity of the linear coefficients and the Weierstrass degrees,
@@
are vertices of the Newton polygon of $f^{\circ n+1}$. By varying $\pi$ among the $p$ roots of $f$, the equations $f^{\circ n}-\pi=0$ describe all the roots of $f^{\circ n+1}$. When $\pi=0$, one then retrieves the roots of $f^{\circ n}$. Otherwise, the Newton polygon of $f^{\circ n}-\pi$ consists of a single line of slope $e/(p^{n+1}-p^{n})$, which is greater than those of the Newton polygon of $f^{\circ n}$ and hence contributes to the last segment. Packing the roots of $f^{\circ n+1}/f^{\circ n}$, that is, the roots of $f^{\circ n+1}$ that are not roots of $f^{\circ n}$, one thus finds that there are $p^{n}(p-1)$ of them and that the Weierstrass polynomial associated to this segment is irreducible over $K$ since $(e,p^{2}-p)=1$. The left-hand side part of the Newton polygon of $f^{\circ n+1}$ comes from the induction hypothesis ∎
@@
    Unlike Specter’s case, the Weierstrass polynomial of $f^{\circ n+1}/f^{\circ n}$ is not Eisenstein in general but still irreducible over $K$, which is sufficient: therefore, the absolute Galois group $G_{K}$ of $K$ acts transitively on the roots of $f^{\circ n+1}/f^{\circ n}$ given that $f$ is defined over $\mathcal{O}_{K}$.
@@
    Asking for the simplicity of the roots of $f$ and its iterates is equivalent to asking that the p1p-1 nonzero roots of $f$ have the same valuation, under the coprimality assumption. This condition is easier to check in practice because it only requires to draw a finite part of the Newton polygon of $f$.
@@
For each nonzero $m\in\mathbb{Z}_{p}$, the fixed points of $u^{\circ m}\!$ are the roots of $f^{\circ v(m)}$ where $v(m)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\upsilon_{K}(u^{\prime}(0)^{m}-1)/e$.
@@
The series $u$ commutes with the noninvertible $f$ having finite Weierstrass degree, ensuring that all of its nontrivial iterates have finitely many fixed points [Lub94, Corollary 4.3.1]. The decreasing part of the Newton polygon of $u^{\circ m}(X)-X$ evolves between the vertex $(1,\upsilon_{K}(u^{\prime}(0)^{m}-1))$ and the $X$-axis. Since $u$ is defined over $\mathcal{O}_{K}$, its fixed points are a finite union of $G_{K}$-orbits in $\Lambda(f)$. According to Remark 1.2, each segment in the latter polygon then makes a decrease of $e$ along the $Y$-axis. ∎
@@ 2. The latent Tate module
As stated in the introduction, the $p$-adic Tate module of a formal group characterizes it up to isomorphism. Specter provided a method to recover the associated representation in [Spe18]. In this section, we adapt this method to our case, and go one step further by retrieving the full structure on the set of $f$-consistent sequences.
@@ 2.1. The Galois action on $f$-consistent sequences
The full structure on $T_{f}$ can be recovered solely from its Galois action. To this end, we are interested in finding a more explicit description of this action. This can be achieved by comparing it with the action of the $\mathbb{Z}_{p}$-iterates of $u$ on $T_{f}$. Recall that $T_{f}$ is the $G_{K}$-set of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}}$, and that the pair of series $(f,u)$ is a pair of endomorphisms of it, acting by componentwise evaluation:
@@
Let $S_{f}\subset T_{f}$ be the subset composed of all elements whose second entry is nonzero. If Lubin’s conjecture holds, then $T_{f}$ is canonically isomorphic to the $p$-adic Tate module of the latent formal group, and an element $x\in T_{f}$ is a generator of $T_{f}$ as $\mathbb{Z}_{p}$-module if and only if $x\in S_{f}$. Elements of $T_{f}$ are either zero or have finitely many initial zero entries followed by an element of $S_{f}$. In order to describe the $G_{K}$-action on $T_{f}$, it is therefore sufficient to describe it on the subset $S_{f}$. However, recovering this action dynamically via the $\mathbb{Z}_{p}$-iterates of $u$ requires restricting it. Indeed, $u$, and hence its iterates, fix some initial entries of all the elements of $T_{f}$. We let
. . $\ell\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\upsilon_{K}(u^{\prime}(0)-1)/e\geqslant 1$
which is a positive integer since we assume $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. According to Corollary 1.3, $u$ fixes up to the $\ell+1$ first components of the elements of $S_{f}$. We view $S_{f}$ as a directed tree:
@@
where the diagram is given in the case $p=3$ and $\ell=2$. Indexed from bottom to top, its vertices at level $n\geqslant 1$ are the roots of $f^{\circ n}/f^{\circ n-1}$, and there is an edge from a vertex $v_{2}$ to a vertex $v_{1}$ if and only if $f(v_{2})=v_{1}$. Given $\pi=(\pi_{n})_{n\geqslant 0}\in S_{f}$, we let
$S_{f,\pi}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{(x_{n})_{n\geqslant 0}\in T_{f}\mid x_{0}=\pi_{0},\dots,x_{\ell}=\pi_{\ell}\}$
which is a subset of $S_{f}$. We set $j\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p^{\ell}-p^{\ell-1}$; then any sequence $(\alpha_{1},\dots,\alpha_{j})$ of elements $\alpha_{i}=(\alpha_{i,n})_{n\geqslant 0}$ of $S_{f}$ such that $\{\alpha_{1,\ell},\dots,\alpha_{j,\ell}\}$ is the set of roots of $f^{\circ\ell}/f^{\circ\ell-1}$ determines a partition $(S_{f,\alpha_{1}},\dots,S_{f,\alpha_{j}})$ of $S_{f}$. In what follows, we first retrieve the Galois action on such subsets, and then discuss of its global action on $S_{f}$. The next result allows us to pass from one element of $S_{f}$ to another.
@@
For every $n\geqslant 1$, the $n$-th component of an element of $S_{f}$ is a root of $f^{\circ n}$ that is not a root of $f^{\circ n-1}$. These roots are the roots of an irreducible polynomial over $K$, as stated in Remark 1.2. Therefore $G_{K}$ acts transitively on this set and, since $f$ is defined over $\mathcal{O}_{K}$, this action passes to the limit as a transitive action. ∎
@@ 2.1.1. Action on cells
Let $\pi=(\pi_{n})_{n\geqslant 0}$ be an element of $S_{f}$, and set $L_{\pi}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=K(\pi_{\ell})$. Then the Galois group $G_{L_{\pi}}\!$ fixes the $\ell+1$ first components of $\pi$. We begin by recovering the action of $G_{L_{\pi}}\!$ on the cell $S_{f,\pi}\subset S_{f}$.
@@
This is the same as [Spe18, Proposition 2.9] adapting the notations; the key point lies in Corollary 1.3. For the reader’s convenience and due to the stature of this result, we outline the proof. To show that the action is transitive, it suffices to show that it is the case on
$O_{n}(\pi_{\ell})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid f^{\circ n-\ell}(x)=\pi_{\ell}\}\subset\Lambda(f)$
for all integers $n>\ell$. Using the Newton polygon, one finds that the cardinality of $O_{n}(\pi_{\ell})$ is $p^{n-\ell}$. Joint to Corollary 1.3, the stabilizer of an element $x\in O_{n}(\pi_{\ell})$ under the action of the $\mathbb{Z}_{p}$-iterates of $u$ is
$\operatorname{Stab}(x)=\big\{u^{\circ m}\in u^{\circ\mathbb{Z}_{p}}\;\big|\;\upsilon_{K}(u^{\prime}(0)^{m}-1)\geqslant ne\big\}=u^{\circ p^{n-\ell}\mathbb{Z}_{p}}$
since we have picked $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. By the orbit-stabilizer theorem, we obtain that $u^{\circ\mathbb{Z}_{p}}$ acts transitively on $O_{n}(\pi_{\ell})$. Consequently, the action of the statement is also free. ∎
The transitive action of $G_{K}$ on the set $S_{f}$ restricts to a transitive action of $G_{L_{\pi}}\!$ on the subset $S_{f,\pi}$. Combining this observation with Proposition 2.2 enables us to recover the latter action via the following rule: as $g$ varies in $G_{L_{\pi}}$, the elements $g.\pi$ exhaust the set $S_{f,\pi}$, and comparing $g.\pi$ with $\pi$ gives a unique $u_{g}\in u^{\circ\mathbb{Z}_{p}}$ satisfying
@@
Using the previous notation, the map $\chi_{f,\pi}\colon G_{L_{\pi}\!}\rightarrow\mathbb{Z}_{p}^{\times}\colon g\mapsto u_{g}^{\prime}(0)$ is a character of $G_{L_{\pi}}\!$ satisfying $g.\pi=[\chi_{f,\pi}(g)]_{f}(\pi)$ for every $g\in G_{L_{\pi}}$.
@@ 2.1.2. Action on the whole set
Having analyzed the situation on cells, we turn to the global case. To compare the local actions with each other, we consider the finite extension $L$ of $K$ generated by the roots of $f^{\circ\ell}/f^{\circ\ell-1}$, or equivalently, the Galois closure of $L_{\pi}/K$ for any $\pi\in S_{f}$. Given $\pi,\pi^{\prime}\in S_{f}$, Proposition 2.1 guarantees the existence of a $\sigma\in G_{K}$ satisfying $\pi^{\prime}=\sigma.\pi$. Consequently $L_{\pi^{\prime}}=\sigma(L_{\pi})$ and the absolute Galois groups of $L_{\pi}$ and $L_{\pi^{\prime}}$ are $\sigma$-conjugate in $G_{K}$. Since $u$ is defined over $\mathcal{O}_{K}$, the previous rule gives, for every $g\in G_{L}$, that
@@
After the two preparatory subsections that follow, we shall prove that this character is crystalline of weight $1$. Therefore, there exists a formal group over $\mathcal{O}_{L}$ whose associated character is $\chi_{f}$, and we transport the structure of its Tate module to $T_{f}$. Theorem 3.1 will ultimately imply that the choice of $\pi$ was independent of the subsequent constructions.

### 22 Apr 2026 v2

@@ 1.2. Torsion points
Whenever we talk about the roots or the fixed points of a power series, we mean its roots or its fixed points, respectively, in the $p$-adic open unit disk $\mathfrak{m}_{\mathbb{C}_{p}\!}$. A noninvertible power series $s_{1}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s_{1}^{\prime}(0)$ is nonzero can have no other fixed points than 0, but many roots. The set of roots of $s_{1}$ and its iterates is denoted by
$\Lambda(s_{1})=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid\exists\,n\geqslant 0,\;s_{1}^{\circ n}(x)=0\}.$
On the other hand, a nontorsion invertible power series $s_{2}\in\mathscr{S}_{0}(\mathcal{O}_{K})$ can have no other roots than 0, but many fixed points. If $s_{1}$ and $s_{2}$ commute with each other, Lubin showed in [Lub94, Proposition 3.2] that $\Lambda(s_{1})$ is the set of periodic points of $s_{2}$. Knowing more about these points (their valuation, how many there are for a given valuation) is a key point in our arguments. This can be achieved by means of the Weierstrass preparation theorem and Newton polygons. Working with Newton polygons requires fixing a valuation; for this, we always use $\upsilon_{K}$ the one normalized on $K$ so that $\upsilon_{K}(p)=e$.
@@
We proceed by induction on $n$. The base case ($n=1$) follows directly from the hypotheses on $f$ and [Li96, Theorem 4.1]. The decreasing part of the Newton polygon of $f$ consists of a single line of slope e/(p1)-e/(p-1), and this implies that the Weierstrass polynomial associated to $f(X)/X$ is irreducible over $K$ given that we assume (e,p1)=1(e,p-1)=1. Assume now that this holds for $n\geqslant 1$. By multiplicativity of the linear coefficients and the Weierstrass degrees,
@@
are vertices of the Newton polygon of $f^{\circ n+1}$. By varying $\pi$ among the $p$ roots of $f$, the equations $f^{\circ n}-\pi=0$ describe all the roots of $f^{\circ n+1}$. When $\pi=0$, one then retrieves the roots of $f^{\circ n}$. Otherwise, the Newton polygon of $f^{\circ n}-\pi$ consists of a single line of slope $e/(p^{n+1}-p^{n})$, which is greater than all those of the Newton polygon of $f^{\circ n}$ and therefore contributes to the last segment. Packing the roots of $f^{\circ n+1}/f^{\circ n}$, that is, the roots of $f^{\circ n+1}$ that are not roots of $f^{\circ n}$, one thus finds that there are $p^{n}(p-1)$ of them and that the Weierstrass polynomial associated to this segment is irreducible over $K$ since we assume $(e,p^{2}-p)=1$. The left-hand side part of the Newton polygon of $f^{\circ n+1}$ comes from the induction hypothesis. ∎
@@
    Unlike Specter’s case, the Weierstrass polynomial of $f^{\circ n+1}/f^{\circ n}$ is not Eisenstein in general but still irreducible over $K$, which is sufficient: therefore the absolute Galois group $G_{K}$ of $K$ acts transitively on the roots of $f^{\circ n+1}/f^{\circ n}$ since $f$ is defined over $\mathcal{O}_{K}$.
@@
    Asking for the simplicity of the roots of $f$ and its iterates is equivalent to asking that the p1p-1 nonzero roots of $f$ have the same valuation, under the coprimality assumption. This reformulation is easier to check in practice because it only requires to draw a finite part of the Newton polygon of $f$.
@@
For every nonzero $m\in\mathbb{Z}_{p}$, the fixed points of $u^{\circ m}\!$ are the roots of $f^{\circ v(m)}$ where $v(m)=\upsilon_{K}(u^{\prime}(0)^{m}-1)/e$.
@@
The series $u$ commutes with the noninvertible $f$ having finite Weierstrass degree, ensuring that its nontrivial iterates have finitely many fixed points by [Lub94, Corollary 4.3.1]. The decreasing part of the Newton polygon of $u^{\circ m}(X)-X$ evolves between the vertex $(1,\upsilon_{K}(u^{\prime}(0)^{m}-1))$ and the $X$-axis. Since $u$ is defined over $\mathcal{O}_{K}$, its fixed points are a finite union of $G_{K}$-orbits in $\Lambda(f)$. According to Remark 1.2, each segment in the latter polygon then makes a decrease of $e$ along the $Y$-axis. ∎
@@ 2. The latent Tate module
As stated in the introduction, the $p$-adic Tate module of a formal group characterizes it up to isomorphism. Specter provided a method in [Spe18] to recover the associated representation. In this section, we adapt this method to our case and go further by retrieving the full structure on the set of $f$-consistent sequences.
@@ 2.1. The Galois action on $f$-consistent sequences
The full structure on $T_{f}$ can be recovered solely from how Galois acts on it. To this end, we are interested in finding a more explicit description of this action. This can be achieved by comparing it with the action of the $\mathbb{Z}_{p}$-iterates of $u$ on $T_{f}$. We recall that $T_{f}$ is the $G_{K}$-set of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}\!}$, and that the pair of series $(f,u)$ is a pair of endomorphisms of it, acting by componentwise evaluation:
@@
Let $S_{f}\subset T_{f}$ be the subset composed of all elements whose second entry is nonzero. If Lubin’s conjecture holds, $T_{f}$ is canonically isomorphic to the $p$-adic Tate module of the latent formal group, and an element $x\in T_{f}$ is a generator of $T_{f}$ as $\mathbb{Z}_{p}$-module if and only if $x\in S_{f}$. Elements of $T_{f}$ are either zero or have finitely many initial zero entries followed by an element of $S_{f}$. In order to describe the $G_{K}$-action on $T_{f}$, it is therefore sufficient to describe it on the subset $S_{f}$. Nevertheless, recovering this action dynamically via the $\mathbb{Z}_{p}$-iterates of $u$ requires restricting it. Indeed, $u$, and hence all its iterates, fix some initial entries of all the elements of $T_{f}$. We let
$\ell=\upsilon_{K}(u^{\prime}(0)-1)/e\geqslant 1$
which is a positive integer since we have picked $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. By Corollary 1.3, $u$ fixes up to the $\ell+1$ first components of the elements of $S_{f}$. We interpret $S_{f}$ as a directed tree, as shown in Figure 1. Indexed from bottom to top, its vertices at level $n\geqslant 1$ are the roots of $f^{\circ n}/f^{\circ n-1}$, and there is an edge from a vertex $v_{2}$ to a vertex $v_{1}$ if and only if $f(v_{2})=v_{1}$. Given $\pi=(\pi_{n})_{n\geqslant 0}\in S_{f}$, we let
$S_{f,\pi}=\{(x_{n})_{n\geqslant 0}\in T_{f}\mid x_{0}=\pi_{0},\dots,x_{\ell}=\pi_{\ell}\}$
which is a subset of $S_{f}$. Set $j=p^{\ell}-p^{\ell-1}$; any sequence $(\alpha_{1},\dots,\alpha_{j})$ of elements $\alpha_{i}=(\alpha_{i,n})_{n\geqslant 0}$ of $S_{f}$ such that $\{\alpha_{1,\ell},\dots,\alpha_{j,\ell}\}$ is the set of roots of $f^{\circ\ell}/f^{\circ\ell-1}$ then determines a partition $(S_{f,\alpha_{1}},\dots,S_{f,\alpha_{j}})$ of $S_{f}$. In what follows, we first retrieve the Galois action on such subsets and then discuss of its global action on $S_{f}$. The next result allows us to pass from one element of $S_{f}$ to another.
@@
Figure 1. Directed tree associated to $S_{f}$ in the case $p=3$ and $\ell=2$.
@@
For each $n\geqslant 1$, the $n$-th component of an element of $S_{f}$ is a root of $f^{\circ n}$ that is not a root of $f^{\circ n-1}$. These roots are the roots of an irreducible polynomial over $K$, as stated in Remark 1.2. Therefore $G_{K}$ acts transitively on this set. To prove that the transitivity passes to the limit, pick $(x_{n})_{n\geqslant 0}$, $(y_{n})_{n\geqslant 0}\in S_{f}$ and define for all $n\geqslant 0$ the set
$A_{n}=\{\sigma\in G_{K}\mid\sigma.x_{n}=y_{n}\}$
which is closed. Since $G_{K}$ is compact, the nested sequence $(A_{n})_{n\geqslant 0}$ has a nontrivial intersection. This proves the proposition. ∎
@@ 2.1.1. Action on cells
Let $\pi=(\pi_{n})_{n\geqslant 0}$ be an element of $S_{f}$ and set $L_{\pi}=K(\pi_{\ell})$. Then the group $G_{L_{\pi}}\!$ fixes the $\ell+1$ first components of $\pi$. We begin by recovering the action of $G_{L_{\pi}}\!$ on the cell $S_{f,\pi}\subset S_{f}$.
@@
This is the same as [Spe18, Proposition 2.9] adapting the notations. For the reader’s convenience and due to the stature of this result, we outline the proof. To show that the action is transitive, it suffices to show, by compacity of $\mathbb{Z}_{p}$, that it is the case on
$O_{n}(\pi_{\ell})=\{x\in\mathfrak{m}_{\mathbb{C}_{p}}\!\mid f^{\circ n-\ell}(x)=\pi_{\ell}\}\subset\Lambda(f)$
for every integer $n>\ell$. Using the Newton polygon, one finds that the cardinality of $O_{n}(\pi_{\ell})$ is $p^{n-\ell}$. Joint to Corollary 1.3, the stabilizer of an element $x\in O_{n}(\pi_{\ell})$ under the action of the $\mathbb{Z}_{p}$-iterates of $u$ is
$\operatorname{Stab}(x)$ $=\big\{u^{\circ m}\in u^{\circ\mathbb{Z}_{p}}\;\big|\;\upsilon_{K}(u^{\prime}(0)^{m}-1)\geqslant ne\big\}$
$=\big\{u^{\circ m}\in u^{\circ\mathbb{Z}_{p}}\;\big|\;\upsilon_{K}(u^{\prime}(0)-1)+\upsilon_{K}(m)\geqslant ne\big\}$
$=u^{\circ p^{n-\ell}\mathbb{Z}_{p}}$
since we have picked $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. By the orbit-stabilizer theorem, we deduce that $u^{\circ\mathbb{Z}_{p}}$ acts transitively on $O_{n}(\pi_{\ell})$. Therefore $u^{\circ\mathbb{Z}_{p}}$ acts transitively on $S_{f,\pi}$, and this action is free because of [Lub94, Corollary 4.3.1]. ∎
The transitive action of $G_{K}$ on the set $S_{f}$ restricts to a transitive action of $G_{L_{\pi}\!}$ on the subset $S_{f,\pi}$. Combining this observation with Proposition 2.2 enables us to recover the latter action via the following rule: as $g$ varies in $G_{L_{\pi}}$, the elements $g.\pi$ exhaust the set $S_{f,\pi}$, and comparing $g.\pi$ with $\pi$ gives a unique $\mathbb{Z}_{p}$-iterate $u_{g}$ of $u$ such that
@@
The map $\chi_{f,\pi}\colon G_{L_{\pi}\!}\rightarrow\mathbb{Z}_{p}^{\times}$ defined by $g\mapsto u_{g}^{\prime}(0)$ is a character of $G_{L_{\pi}}\!$ satisfying $g.\pi=[\chi_{f,\pi}(g)]_{f}(\pi)$ for every $g\in G_{L_{\pi}\!}$.
@@ 2.1.2. Action on the whole set
Having analyzed the situation on cells, we turn to the global case. To compare the local actions with each other, we consider the finite extension $L$ of $K$ generated by the roots of $f^{\circ\ell}/f^{\circ\ell-1}$ or, equivalently, the Galois closure of $L_{\pi}/K$ for any $\pi\in S_{f}$. Given $\pi,\pi^{\prime}\in S_{f}$, Proposition 2.1 guarantees the existence of a $\sigma\in G_{K}$ satisfying $\pi^{\prime}=\sigma.\pi$. Therefore $L_{\pi^{\prime}}=\sigma(L_{\pi})$ and the absolute Galois groups of $L_{\pi}$ and $L_{\pi^{\prime}}$ are $\sigma$-conjugate in $G_{K}$. Since $u$ is defined over $\mathcal{O}_{K}$, the previous rule gives, for every $g\in G_{L}$, that
@@
After the two preparatory subsections that follow, we shall prove that this character is crystalline of weight $1$. Consequently there exists a formal group defined over $\mathcal{O}_{L}$ whose associated character is $\chi_{f}$, and we transport the structure of its $p$-adic Tate module to $T_{f}$. Theorem 3.1 will ultimately imply that the subsequent constructions were independent of the choice of $\pi$.